OpenStudy (anonymous):

David has a hose that is 27 ft long. He wants to cut it into two pieces so that the longer piece is 3 ft more than 2 times the length of the shorter piece. How long will each piece of hose be? Shorter piece: ft Longer piece: ft

OpenStudy (littlebird):

You should make a formula

OpenStudy (littlebird):

Let's say x is the length of the shorter piece. Since the rope is 27 ft long, I will set the equation equal to 27. It will look like this... short piece + long piece= 27 the long piece is 3 ft more than two times the short piece (2x+3) So all together I have... x+(2x+3)=27

OpenStudy (anonymous):

Is it Shorter: 12ft Longer: 15ft

OpenStudy (anonymous):

lol it would not let me post so i was like typing forever

OpenStudy (anonymous):

it look like

OpenStudy (littlebird):

Add like terms 3x+3=27 Solve for x 3x=24 X=8 The shorter piece is 8 The long piece is 2(8)+3 which is 19 short= 8 long=19 Since 19+8=27 I think my answer is correct

OpenStudy (anonymous):

oh ya it was right

OpenStudy (anonymous):

sorry here is a medal

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

thank you @Littlebird